OpenStudy (anonymous):

I need to leave the secx alone Sec^2x-1=secx+1 I need to know who answer anothers comments sorry i am new so i don't nothing!!

OpenStudy (akashdeepdeb):

No problem! We are all here to help. :) What do you need to solve for in this question?

OpenStudy (akashdeepdeb):

I think, what you mean by leave sec x alone is that we have to solve for it. Now let sec x = y You are given: sec^2 x - 1 = sec x + 1 y^2 - y - 2 = 0 Factorize. y^2 - 2y + y - 2 = 0 y(y-2) + 1(y-2) = 0 (y+1) (y-2) = 0 So, basically: y = -1,2 Which implies that: y = sec x = -1,2. Getting this? :)